# CompilerDream: Learning a Compiler World Model for General Code Optimization

### v2

@@ Abstract
Effective code optimization in compilers is crucial for computer and software engineering. The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code. While most compilers rely on a fixed sequence of optimization passes, current methods to find the optimal sequence either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training. We introduce CompilerDream, a model-based reinforcement learning approach to general code optimization. CompilerDream comprises a compiler world model that accurately simulates the intrinsic properties of optimization passes and an agent trained on this model to produce effective optimization strategies. By training on a large-scale program dataset, CompilerDream is equipped to serve as a general code optimizer across various application scenarios and source-code languages. Our extensive experiments first highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard. More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and other state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
1 Introduction
@@
Code optimization plays an important role in realizing the full potential of software and hardware. Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort. To achieve this, compilers usually employ a front-end that translates the source code of any language into an intermediate representation (IR), a back-end that converts IR to the binary code optimizer, and a middle-end optimizer that performs language- and platform-agnostic optimizations over the IR of the original code (Figure 1). The optimizer is typically implemented as a series of passes applying transforms on the code. The performance of the optimization primarily depends on the selection and order of the optimization passes. Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance. For example, -O1, -O2, and -O3 improve for execution speed, and -Os and -Oz for reducing program size. Obviously, given the vast diversity of programs and platforms, these off-the-shelf optimization strategies predefined by compiler experts are not optimal for most circumstances, and we can significantly improve the performance of a specific program if we can find a better optimization pass sequence compared to that provided by compilers (Triantafyllis et al. 2003; Georgiou et al. 2018).
Figure 1: CompilerDream agent performs code optimization using a learned policy in a compiler environment. Guided by a learned policy, at each step, CompilerDream analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics. Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
@@
Automatically optimizing the pass sequence is thus crucial in improving the compiler optimization performance. To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs. However, current research often fails to meet these requirements simultaneously. Search-based methods (Bodin et al. 1998) ensure near-optimality but require thousands of compilations for each new program, making it too time-consuming for regular use. In contrast, machine learning methods can either predict the optimization sequences directly or estimate the outcomes of applying an optimization sequence to guide a search, thereby avoiding the time-consuming process of interacting with the compiler to validate the optimization outcome. However, in addition to potentially sacrificing some optimality compared to searching, these learning-based methods face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
@@
A range of machine learning applications (Devlin et al. 2019; Brown et al. 2020; Radford et al. 2021; Kirillov et al. 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al. 2022). However, the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al. 2020; Shahzad et al. 2022) or from relatively small training sets comprising only a few hundred programs (Mammadli, Jannesari, and Wolf 2020; Jain et al. 2022). Additionally, the predominant use of model-free reinforcement learning methods with limited capacity may also hinder generalization.
In this study, we focus on the LLVM (Lattner and Adve 2004) phase ordering problem, a longstanding challenge for compiler research, and propose CompilerDream, a world model-based reinforcement learning approach capable of generating an optimization pass sequence tailored to any particular input program. Our model-based method (Hafner et al. 2023) not only learns an accurate predictive world model of compilers to simulate real compiler executions, whose number can be greatly reduced, but also benefits generalization by learning representations that better capture the structure of the compiler state transitions (Anand et al. 2022). Trained on a carefully curated large-scale dataset of natural programs, the world model accurately captures the dynamics of the compiler optimization process, enabling the optimization agent to generalize to unseen programs and predict superior optimization sequences.
@@
We demonstrate the effectiveness of CompilerDream on a range of domains (Cummins et al. 2022), from benchmark suites encompassing fundamental algorithms to production-level open-source programs, including object files from C++ TensorFlow (Abadi et al. 2016) and OpenCV (Culjak et al. 2012) library. Even without considering generalization, CompilerDream’s strong optimization capabilities place it at the top of the CompilerGym leaderboard for autotuning. For general code optimization, the large-scale trained world models of CompilerDream can accurately predict the outcomes of pass sequences on unseen programs, a setting known as value prediction. Beyond selecting optimization sequences based on value prediction, the agent can also directly produce optimization sequences in a single trial, yielding more efficient code size reduction compared to both the -Oz flag and state-of-the-art methods.
The main contributions of this work are three-fold:
@@
  • We introduce CompilerDream, the first model-based reinforcement learning approach for code optimization.
@@
  • We first leverage a large-scale dataset, CodeContests, to achieve strong generalization on unseen programs.
  • Extensive experiments spanning autotuning, value prediction, and end-to-end code optimization, demonstrate the superior performance of CompilerDream.
2 Related Work
@@
One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation (Bodin et al. 1998) or autotuning (Datta et al. 2008). However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy. This limitation underscores the importance of integrating machine learning techniques.
Supervised Learning
@@
Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins 2020). The first approach requires an extensive search on each training programs to identify the most effective optimization sequence, which then serves as the data labels. An early example (Calder et al. 1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al. 2008), a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one (Stephenson et al. 2003; Luk, Hong, and Kim 2009).
Reinforcement Learning
@@
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data (Kulkarni and Cavazos 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al. 2021), loop transformation (Haj-Ali et al. 2020; Brauckmann, Goens, and Castrillón 2021), and graph partitioning (Mirhoseini et al. 2017). Several works relevant to us have explored the full optimization pipeline, i.e. the LLVM phase ordering problem, including AutoPhase (Haj-Ali et al. 2020), CORL (Mammadli, Jannesari, and Wolf 2020), and POSET-RL (Jain et al. 2022).
Model-based Reinforcement Learning
While existing work on reinforcement learning for code optimization predominantly focuses on model-free RL methods, our work pioneers the use of an advanced model-based RL approach. Model-based RL (Sutton and Barto 2018) can offer advantages in terms of both sample efficiency and generalization. Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming. This is further compounded when constructing complex observations, such as control-data flow graphs (Brauckmann et al. 2020; Cummins et al. 2021). Model-based RL addresses these challenges by learning a world model that approximates state transitions and reward signals of the compiler. This is related in spirit to the aforementioned approach of supervised performance models but goes further by also learning a policy through simulating trajectories based on model predictions. This approach thus improves sample efficiency not only by eliminating the necessity for a guided search process during deployment but also by reducing trial-and-error interactions with the real compiler environment during training. Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy (Anand et al. 2022; Mazoure et al. 2022; Ma et al. 2024).
Figure 2: Design overview of CompilerDream. (a) In the model-based training loop, the CompilerDream agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations. (b) The trained agent can be deployed and is capable of zero-shot generalization to unseen programs, delivering effective optimizations. (c) The compiler environment is set up with the input program’s IR, $s_{0}$. At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements. (d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal neural states, and predicting future observations and rewards in response to input actions. (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
3 Method
@@
In this section, we first present the mathematical formulation of the code optimization problem (Section 3.1). Then we introduce the CompilerDream approach for both compiler world model learning (Section 3.2) and optimization agent learning (Section 3.3). Finally, we elaborate on considerations of curating large-scale training dataset (Section 3.4).
3.1 Phase Ordering Decision Process
@@
As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, which is also referred to as the phase ordering problem. It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto 2018). The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space. Complex observation spaces offer comprehensive program information, while expert-designed features (Haj-Ali et al. 2020) embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
A code optimization agent guided by policy $\pi$ determines the optimization sequence of an input program through a series of interactions with the compiler environment. At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$ and its policy $\pi(a_{t}\mid o_{\leq t})$, which transforms the current state $s_{t}$ to the next state $s_{t+1}=p(s_{t},a_{t})$ and will receive a reward $r_{t}$. The goal of the agent is to learn an optimization policy $\pi$ that effectively maximizes the cumulative rewards $G=\sum_{t}r_{t}$, thereby optimizing the performance metrics of the final IR.
3.2 Learning a Compiler World Model

### v3

@@ Abstract.
Effective code optimization in compilers is crucial for computer and software engineering. The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code. While most compilers rely on fixed pass sequences, current methods to find the optimal sequence for specific programs either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training. To address these challenges, we introduce CompilerDream, the first world-model-based approach for general code optimization. CompilerDream features a compiler world model with a reward smoothing technique, enabling accurate simulation of optimization processes. Built on this model, code optimization agents can then be constructed via value prediction or direct optimization sequence generation. Trained on a large-scale program dataset, these agents serve as versatile code optimizers across diverse application scenarios and source-code languages. Our extensive experiments highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard. More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and state-of-the-art methods in both settings of value prediction and end-to-end code optimization. [^1] Code and data available at https://github.com/thuml/CompilerDream.
World Models, Compiler Optimization, Code Data Mining
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736887isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learningccs: Software and its engineering Compilersccs: Computing methodologies Model development and analysis
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15552746. The data of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15549673.
Figure 1. CompilerDream performs code optimization by interacting with the compiler using an agent powered by a world model. With strong generalization capabilities, it efficiently optimizes programs from diverse origins.
[…]
Diverse source code, including C++, C, Rust, and other languages, is first translated into an intermediate representation (IR) by the compiler frontend. CompilerDream, leveraging a world model, optimizes the IR by selecting an optimal pass at each step using an optimization agent. The optimized IR is then processed by the compiler backend to generate a high-quality executable file.
1. Introduction
@@
Code optimization plays an important role in realizing the full potential of software and hardware. Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort. Compilers achieve this through a front-end that translates source code into an intermediate representation (IR), a middle-end optimizer that performs language- and platform-agnostic IR optimizations, and a back-end that converts IR to the binary code (Figure 1). The optimizer is typically implemented as a series of passes applying transforms on the code, where performance largely depends on the selection and order of these optimization passes. Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance, such as -O1, -O2, and -O3 for execution speed, and -Os and -Oz for program size reduction.
@@
Obviously, given the vast diversity of programs and platforms, these off-the-shelf strategies predefined by compiler experts are suboptimal for most circumstances. Automatically optimizing the pass sequence for specific programs thus can yield significant performance gains over default compiler settings (Triantafyllis et al., 2003; Georgiou et al., 2018). To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs. However, current research often fails to meet these requirements simultaneously. Search-based methods (Bodin et al., 1998) achieve near-optimality but require thousands of compilations per program to validate the optimization outcomes, making them impractical. In contrast, machine learning methods avoid these time-consuming compiler interactions, by either predicting the optimization sequences directly or estimating the outcomes of optimization sequences to guide a search.
@@
However, these learning-based methods still risk sacrificing some optimality and, more importantly, face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples. A range of machine learning applications (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021; Kirillov et al., 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
This paper focuses on the LLVM (Lattner and Adve, 2004) phase ordering problem, a longstanding challenge for compiler research. We propose CompilerDream, a world-model-based approach for general code optimization, capable of handling a wide variety of programs, unlike most prior approaches focusing on narrow, domain-specific ones. It learns an accurate predictive world model to simulate compiler executions, forecasting future IR states and optimization metric improvements based on the current IR state and applied pass. We believe that employing a world model offers the following advantages for compiler optimization: First, by capturing optimization dynamics, the world model gains generalizable knowledge into our method (Anand et al., 2022), improving the overall generalization performance. Second, it replaces costly compiler invocations, significantly reducing computational overheads of search and learning algorithms and facilitating large-scale training. Additionally, its high-capacity architecture can extract deeper insights from extensive datasets.
Specifically, to better adapt the world model for compiler optimization, we closely examine the optimization process and introduce a reward smoothing technique to enhance world model training and improve simulation accuracy. Built on this accurate world model, CompilerDream supports various types of optimization agents, including value prediction and reinforcement learning. Moreover, we carefully curate a large-scale training dataset of natural programs, enhancing the world model’s and agent’s generalization to unseen programs, enabling our method to predict superior optimization sequences.
@@
We demonstrate CompilerDream’s effectiveness across a range of program domains (Cummins et al., 2022) and problem scenarios. Our test domains include benchmark suites covering fundamental algorithms, as well as production-level open-source programs, such as object files from C++ TensorFlow (Abadi et al., 2016) and OpenCV (Culjak et al., 2012) libraries. Even without considering generalization, CompilerDream’s strong optimization capabilities top the CompilerGym leaderboard for autotuning. For general code optimization, our large-scale trained world model accurately predicts the outcomes of pass sequences on unseen programs. Both types of optimization agents equipped with the world model outperform the built-in -Oz flag and state-of-the-art methods in their respective settings. In the value prediction setting, the world model serves as an accurate value predictor, enabling superior action sequence selection. In the reinforcement learning setting, the agent trained entirely in the world model directly generates optimization sequences in a single trial, achieving more efficient code across diverse datasets.
The main contributions of this work are as follows:
@@
  • We propose CompilerDream, the first world-model-based approach for code optimization.
  • We introduce a reward smoothing technique that facilitates the application of world models to compiler optimization, benefiting future research in this area.
  • Our approach supports multiple optimization agents, including value prediction and reinforcement learning agents, unifying different methods in this field.
@@
  • By leveraging the large-scale CodeContests dataset and the world model’s generalization capability, we achieve strong performance on diverse unseen programs.
  • Extensive experiments across various program domains and problem scenarios, including autotuning, value prediction, and end-to-end code optimization, show that CompilerDream outperforms state-of-the-art methods.
2. Related Work
@@
A key challenge in compilation is determining which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This involves effectively searching and evaluating numerous options, a process known as iterative compilation (Bodin et al., 1998) or autotuning (Datta et al., 2008). However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy. This limitation underscores the importance of integrating machine learning techniques.
Supervised learning.
@@
Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins, 2020). The first approach requires an extensive search for each training program to identify the most effective optimization sequence, which then serves as the data labels. An early example (Calder et al., 1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al., 2008), a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables the evaluation of possible options without the need to compile and profile each one (Stephenson et al., 2003; Luk et al., 2009). Coreset-NVP (Liang et al., 2023) follows this approach and achieves state-of-the-art performance.
Reinforcement learning.
@@
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the need for collecting optimal labeled data (Kulkarni and Cavazos, 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017). Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
World model.
A world model (Ha and Schmidhuber, 2018; LeCun, 2022) that approximates state transitions and reward signals is typically used in two ways: (1) it enables simulation of ”unseen” interactions that are unavailable or unaffordable (Janner et al., 2019; Hafner et al., 2020); (2) as an auxiliary learning task, it aids in better representation that captures the underlying structure of the environment (Anand et al., 2022; Mazoure et al., 2022; Ma et al., 2024). To the best of our knowledge, we are the first to introduce world models for code optimization. In this context, a world model can function as a compiler simulator, approximating IR transformations and eliminating the need for costly execution and profiling of extensive optimization sequences. Furthermore, by sharing representations with the world model, the policy can generalize more effectively to unseen programs. Thus, model-based agents (Sutton and Barto, 2018) can be implemented, which offer superior sample efficiency and generalization compared to model-free methods.
3. Method
@@
This section first defines the code optimization problem and design choices for observation, action, and reward (Section 3.1). We then present the design and training technique of CompilerDream’s world model (Section 3.2), two optimization agents (Section 3.3), and considerations for large-scale dataset curation (Section 3.4).
3.1. Phase Ordering Decision Process
@@
As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, also known as the phase ordering problem. It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto, 2018). The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space, capturing useful features.
A code optimization agent determines the optimization sequence for an input program through interactions with the compiler environment. At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$, transforming current state to $s_{t+1}=p(s_{t},a_{t})$ and receiving reward $r_{t}$. The agent can employ effective strategies, such as search algorithms and parametric neural networks, to select the best actions, as detailed in Section 3.3.
Under the POMDP formulation, we design the observation and action space as summarized in Table 1. The observation features were selected for efficiency and effectiveness. In preliminary experiments, we find that complex program features like ProGraML (Cummins et al., 2021) and inst2vec (Ben-Nun et al., 2018) significantly slow down CompilerDream with marginal performance gains. In contrast, expert-designed Autophase features (Haj-Ali et al., 2020) leverage domain knowledge, enhancing generalization by filtering irrelevant details. We construct the observation by concatenating the 56-dimensional Autophase feature vector with a 42-dimensional action histogram vector, which records the number of times each action has been selected during the current episode. Both vectors are normalized for consistency: each Autophase feature is divided by the program’s initial total instruction count, and the action histogram is scaled by the per-episode action limit, which is set to 45. We adopt two distinct action spaces in our experiments to align with the baseline methods. The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020) consists of 42 actions. Originally, this action space included 45 LLVM passes, but CompilerGym excludes 3 due to updates in the latest LLVM version, leaving 42 actions. This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020; Cummins et al., 2022).
(1) $r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$
where lower $C$ indicates better performance. Following prior work on code size reduction, we define $C(s)$ as the IR instruction count. $C(s_{\mathrm{b}})$ is the baseline performance achieved by the built-in -Oz flag.
Table 1. Observation and action space in CompilerDream.
Aspect Name Description Dim.
Observation
Autophase
(Haj-Ali et al., 2020)
A feature vector capturing
various IR code statistics.
56
Action
Histogram
A vector where each
dimension represents the
execution frequency of a
specific action.
124
/42
Action
(Sec. 4.2 & 4.3)
Full LLVM
Passes
Full action space with all
LLVM optimization passes.
124
Action
(Sec. 4.4)
Autophase
Passes
A reduced action space
derived from Autophase(Haj-Ali et al., 2020).
42
Figure 2. Design overview of CompilerDream. The world model is trained on real episodes generated by the agent, incorporating a reward smoothing process to stabilize training and enhance learning efficiency. CompilerDream supports training with RL and value prediction agents, which learn from episodes simulated by the world model.
[…]
The agent interacts with a compiler to collect real episodes, which are stored in a replay buffer. While training the world model with these episodes, a reward smoothing process is applied to enhance training stability and learning efficiency. Once trained, the world model functions as a compiler environment simulator, enabling reinforcement learning (RL) agents and value prediction agents to interact with it and generate simulated episodes. These simulated episodes are used to train the agents, ultimately leading to improved performance.
3.2. Learning a Compiler World Model